I am having a problem setting up this problem - Can you help me?
At 8 am, the Smiths left a campground, driving at 48 miles per hour. At 8:20 am, the Garcias left the same campground and followed the same route, driving at 60 miles per hour. At what time did they overtake the Smiths?
I know
rate x time = distance
but that is as far as I got.

Answer by Alan3354(69443)

At 8 am, the Smiths left a campground, driving at 48 miles per hour. At 8:20 am, the Garcias left the same campground and followed the same route, driving at 60 miles per hour. At what time did they overtake the Smiths?
---------------------
20 min = 1/3 hour
In 1/3 hour, the Smiths are 16 miles away (48 mi/hr * 1/3 hr)
The Garcias gain on them at 12 mi/hr (60 - 48)
d = r*t
t = d/r = 16 mi/12 mi/hr = 1 1/3 hours = 1 hr 20 mins
8:20 + 1:20 = 9:40